TeXML

June 23, 1999

What is TeXML?

TeXML is a key component of a path to print solution for XML.

The Path to Print

The path to print begins with your XML document. You write an XSL transform which accepts your document type and outputs a new XML document which conforms to the TeXML document type. The java program, TeXMLatte transforms any document conforming to the TeXML document type into TeX.

You may now use your TeX processor to produce typeset output from XML. The materials provided here give several examples of XML content trans-

formed to TeX.

The Parts of TeXML

There are three parts to the solution:

TeXML.dtd an XML document type definition.

TeXML.java a program written in java which takes as input a DOM conforming to TeXML.dtd and outputs TeX.

TeXMLatte.java a program written in java which takes as input an XML document conforming to TeXML.dtd and outputs TeX. The program parses the document using the XML4J parser. It sends the resulting DOM to a method the TeXML class.